14 divided by 2 7/11
Answer by stanbon(75887)

14 divided by 2 7/11
----------------
14/[(11*2+7)/11]
---
15/[20/11]
---
= (15*11)/20
= (3/4)11
= 33/4
= 8 1/4
============
Cheers,
Stan H.
============